# Low-Dimension-to-High-Dimension Generalization And Its Implications for Length Generalization

### v1

@@ 5.1 Chain-of-Thought for Length Generalization
While the Chain-of-Thought (CoT) can lead to more variety in the length of testing samples, it is widely used as an effective technique to improve the length generalization in various reasoning tasks. This seems contradictory if considered in the original space of the input sequence. However, when considering the abstraction of the data generating process from the hidden space, we can see that CoT intrinsically changes the underlying hidden space by extending each dimension with a “middle” variable and does not lead to the dimensional increase in the hidden space. For example, consider the $n$-addition without CoT and the $n$-addition with CoT. In the case without CoT, the instance $x_{n-1}\dots x_{0}+y_{n-1}\dots y_{0}=z_{n}\dots z_{0}$ corresponds to the latent state $h_{n}=\left[\left(x_{0},y_{0}\right),\dots,(x_{n-1},y_{n-1})\right]\in\Sigma^{n}$ for $\Sigma=\{0,\dots,9\}\times\{0,\dots,9\}$. In the case with CoT, one step of predicting $z_{t}$ in solving the instance corresponds to the latent state $h_{n}=\left[\left(x_{0},y_{0},z_{0}\right),\dots,\left(x_{t-1},y_{t-1},z_{t-1}\right),\left(x_{t},y_{t},*\right)\dots,(x_{n-1},y_{n-1},*)\right]\in\bar{\Sigma}^{n}$ for $\bar{\Sigma}=\Sigma\times\{*,0,\dots,9\}$, where * is a special element indicating undetermined values. CoT does not cause the LDHD generalization challenge but extends the field $\Sigma$, which could potentially lead to a more easily learnable target concept under the model architecture.
@@ 5.2 Position Embeddings for Length Generalization
Position embeddings are considered closely related to the length generalization in Transformers Our analysis suggests a principle to consider both the inherent LDHD generalization and the nuisances such as the data format in the design of the position embeddings. To further elaborate, consider the length generalization of the URF addition with CoT. While RPE can capture the recursive structure of the addition problem and could lead to LDHD generalization in the latent space, it fails to work for the URF addition with CoT. The failure of RPE is attributed to the nuisance of the URF.
@@
$\sum_{1\leq k\leq j,1\leq l\leq i}\frac{\exp\left((W_{Q}x_{j})^{\intercal}(W_{K}x_{l})\right)}{\sum_{1\leq l^{\prime}\leq j}\exp\left((W_{Q}x_{j})^{\intercal}(W_{K}x_{l^{\prime}})\right)}\frac{\exp\left((W_{Q}x_{i})^{\intercal}(W_{K}x_{k})\right)}{\sum_{1\leq k^{\prime}\leq i}\exp\left((W_{Q}x_{i})^{\intercal}(W_{K}x_{k^{\prime}})\right)}R_{(j-l)-(i-k)},$
[…]
Figure 1: Length generalization of Transformer with RPE and RPE-Square in the URF addition. The training data are from the URF $4$-addition and the testing data are from the URF $5$-addition. (a) Both the models are trained for $20000$ steps. The comparison result shows that the RPE fails while the RPE-Square succeeds in achieving length generalization in the URF addition, (b) The RPE-Square model is trained for $200000$ steps, which shows the RPE-Square model converges to a solution achieving nearly perfect accuracy. The “$k$-th” line represents the digitwise accuracy to predict the $k$-th digit, i.e. $z_{k}$.
@@ 6 Conclusion and Discussion
We propose considering LDHD generalization, which captures the inherent scaling challenge of length generalization in various reasoning tasks. We introduce the No-Free-Lunch Theorem for LDHD generalization, demonstrating the necessity of inductive bias for achieving length generalization. We further investigate the inductive biases of different model architectures trained with (S)GD in the context of Boolean functions. LDHD generalization can only be achieved if the inductive bias aligns with the structure of the target concept. As implications for length generalization in practical problems, our perspective on LDHD generalization elucidates the role of CoT in extending the latent space and leads to the principle that both inherent LDHD generalization and nuisances must be addressed to achieve length generalization.
@@
For future work, while our theory is established with simplified models, it is crucial to further investigate inductive bias in more practical and complex models. Additionally, developing a paradigm for position embedding design based on the proposed principles is a valuable avenue for exploration.
Appendix A Proofs
A.1 Proof for Theorem 1
@@
$V_{1}(c):=\sum_{f\in{\mathcal{F}}/\left(c|_{{\mathcal{X}}_{N_{0}}}\right)}\mathbb{E}_{x\sim{\mathcal{D}}_{1}}\left[\ell\left(c(x),f(x)\right)\right],$
$V_{2}(c):=\sum_{f\in{\mathcal{F}}/\left(c|_{{\mathcal{X}}_{N_{0}}}\right)}\mathbb{E}_{x\sim{\mathcal{D}}_{2}}\left[\ell\left(c(x),f(x)\right)\right].$
@@
=$\overset{(a)}{=}\sum_{f\in{\mathcal{F}}/\left(c_{2}|_{{\mathcal{X}}_{N_{0}}}\right)}\mathbb{E}_{x\sim{\mathcal{D}}_{1}}\left[\ell\left(c_{2}(x),c_{2}(x)\right)\right]$
@@
By the no-free-lunch theorem [25, 24] for ${\mathcal{X}}_{N}\setminus{\mathcal{X}}_{N_{0}}$, ${\mathcal{Y}}$, and ${\mathcal{F}}_{{\mathcal{X}}_{N}\setminus{\mathcal{X}}_{N_{0}},{\mathcal{Y}}}$, we have
@@ A.2 Proof for Theorem 2
According to the proof of Theorem 3.8 in Abbe et al. [1], the RFM model $f_{\text{RFM}}(z;a)$ converges to the min-degree interpolator (w.r.t. the variable $z$) in ${\mathcal{G}}^{\prime}_{N_{0},c}$, i.e.,
@@
A.3 Proof for Theorem 5
Since Theorem 3 is a special case of Theorem 5 for ${\mathcal{U}}=\{E_{ij}\}_{1\leq i\leq j\leq N}$, we only need to prove Theorem 5.

### v3

@@ 5.1 Chain-of-Thought for Length Generalization
While the Chain-of-Thought (CoT) can lead to more variety in the length of testing samples, it is widely used as an effective technique to improve the length generalization in various reasoning tasks. This seems contradictory if considered in the original space of the input sequence. However, from the LDHD generalization perspective, we can see that CoT intrinsically changes the underlying hidden space by extending each dimension with a “middle” variable and does not lead to the dimensional increase in the hidden space. For example, consider the $n$-addition without CoT and the $n$-addition with CoT. In the case without CoT, the instance $x_{n-1}\dots x_{0}+y_{n-1}\dots y_{0}=z_{n}\dots z_{0}$ corresponds to the latent state $h_{n}=\left[\left(x_{0},y_{0}\right),\dots,(x_{n-1},y_{n-1})\right]\in\Sigma^{n}$ for $\Sigma=\{0,\dots,9\}^{2}$. In the case with CoT, one step of predicting $z_{t}$ corresponds to the latent state
$\begin{gathered}h_{n}=\left[\left(x_{0},y_{0},z_{0}\right),\dots,\left(x_{t-1},y_{t-1},z_{t-1}\right),\right.\\ \left.\left(x_{t},y_{t},*\right)\dots,(x_{n-1},y_{n-1},*)\right]\in\bar{\Sigma}^{n}\end{gathered}$
for $\bar{\Sigma}=\Sigma\times\{*,0,\dots,9\}$, where * is a special element indicating undetermined values. CoT does not cause the LDHD generalization challenge but extends the domain $\Sigma$, leading to a more easily learnable target concept.
@@ 5.2 Position Embeddings for Length Generalization
Position embeddings are considered closely related to the length generalization in Transformers. Our analysis suggests a principle for position embedding design: consider the inherent LDHD generalization and the data format nuisance separately. To further elaborate, consider the length generalization of the URF addition with CoT. While RPE can capture the recursive structure of the addition problem and could lead to LDHD generalization in the latent space (Zhou et al., 2023; Jelassi et al., 2023), it fails to work for the URF addition with CoT, due to the nuisance of the URF.
@@
$\begin{gathered}\sum_{1\leq k\leq j,1\leq l\leq i}\frac{\exp\left((W_{Q}x_{j})^{\intercal}(W_{K}x_{l})\right)}{\sum_{1\leq l^{\prime}\leq j}\exp\left((W_{Q}x_{j})^{\intercal}(W_{K}x_{l^{\prime}})\right)}\\ \times\frac{\exp\left((W_{Q}x_{i})^{\intercal}(W_{K}x_{k})\right)}{\sum_{1\leq k^{\prime}\leq i}\exp\left((W_{Q}x_{i})^{\intercal}(W_{K}x_{k^{\prime}})\right)}R_{(j-l)-(i-k)},\end{gathered}$
where $W_{Q}$ and $W_{K}$ are the weight matrices for the query and the key, respectively. We replace $\text{RPE}_{j-i}$ with $\text{RPE-Square}_{i,j}$ in the Transformer.
RPE-Square incorporates prior knowledge of LDHD generalization and unaligned data formats by combining RPE with a mechanism to handle unaligned formats. The position embedding for a query $j$ and a key $i$ is determined by the relative distance of relative distances–the difference between the relative distance of $j$ to the position of some token $x_{l}$, and $i$ to the position of some token $x_{k}$, parameterized by $R_{(j-l)-(i-k)}$. This design, which inspires the name RPE-Square, uses a weighted average over all $1\leq l\leq j,1\leq k\leq i$, where the weights are derived from the product of attention scores between $x_{j}$ and $x_{l}$, and $x_{i}$ and $x_{k}$. This approach enables automatic learning of special tokens and is particularly suited for tasks involving unaligned data formats, such as URF addition. Further illustration is provided in Section D.1.
We compare the length generalization performance of RPE and RPE-Square in two tasks: Unaligned Copy and URF Addition. In the unaligned copy, the input is a string whose length is not aligned to a fixed length, and the target is one copy of the input string. An unaligned copy instance of scale $n$ is like $\mathtt{[BOS]\,x_{0}\,\dots\,x_{n-1}\,=\,x_{0}\,\dots\,x_{n-1}\,[EOS]}$”. The URF addition is illustrated in Example 1. To examine length generalization, the models are trained only on small-scale instances but evaluated on instances of larger scales. More details of the experiments are in Section D.2. The experiment results are presented in Figure 2. In the experiments, RPE-Square, the position embedding derived according to the LDHD generalization perspective, can achieve length generalization in both tasks, while RPE fails. The experiment shows that RPE-Square effectively improves over RPE in handling the unalignment in the data format.
@@ 6 Conclusion and Discussion
We propose the LDHD generalization perspective for length generalization, which disentangles the problem into two aspects: LDHD generalization in the latent space and handling data format nuisances. From this perspective, we introduce the No-Free-Lunch Theorem of LDHD generalization, highlighting the necessity of inductive bias for achieving length generalization. Using Boolean functions, we investigate the inductive biases of different model architectures trained with (S)GD and how these inductive biases contribute to LDHD generalization. Our perspective on LDHD generalization further elucidates the role of CoT and leads to the principle of position embedding design in length generalization.
@@
For future work, while our theory is established with simplified models, it is crucial to investigate inductive bias in more practical and complex models. Additionally, developing a paradigm for position embedding design based on the proposed principles is a valuable avenue.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Acknowledgements
Z. Lin and Y. Liang were supported by National Key R&D Program of China (2022ZD0160300). Z. Lin was also supported by the NSF China (No. 62276004). Y. Liang was additionally supported by CCF Baidu Open Fund.
Appendix A Background on Boolean Analysis
In this section, we include a brief background on Boolean analysis essential to this work. We refer to O’Donnell (2014) for further details and comprehensive coverage of Boolean analysis.
Fourier Expansion. A Boolean function $f:\{-1,1\}^{n}\mapsto{\mathbb{R}}$ can always be represented by
$f(x)=\sum_{T\subset[n]}\hat{f}(T)\chi_{T}(x),$ (1)
where $\chi_{T}(x)=\pi_{i\in[T]}x_{i}$. The polynomial in (1) is called the Fourier expansion of the Boolean function $f$. The number $\hat{f}(T)$ is the Fourier coefficient of $f$ on $T$. The set $\{\chi_{T}(x)\}_{T\subset[n]}$ forms a basis, named Fourier basis, for the product space w.r.t the inner product defined by
$\langle f,g\rangle={\mathbb{E}}_{x\sim U\left(\{-1,1\}^{n}\right)}\left[f(x)g(x)\right].$
Degree and Degree Profile. The degree of a Boolean function $f$ is the degree of its Fourier expansion, which is a polynomial. The degree profile of $f$, denoted by $\operatorname{DegP}(f)$, is a $(n+1)$-tuple where $\operatorname{DegP}_{i}(f)=\sum_{T\subset[n],|T|=n+1-i}\hat{f}^{2}(T)$. The order between two degree profiles is lexicographic. The degree profile can be roughly seen as the distribution of the degrees of the monomials in the polynomial. Intuitively, the degree profile reflects the “complexity” of the Boolean function. A lower-degree-profile Boolean function uses fewer variables or combines them more simply than a higher-degree-profile one.
Appendix B Detailed Explanation of PLAA
In this section, we illustrate how PLAA and its variants abstract the effect of position embedding on the LDHD generalization of decoder-only Transformers.
B.1 Construction of PLAA
A typical linear attention at query $n$ can be expressed as
$f_{\text{LA}}\left(x;W_{Q},W_{K},W_{V},B\right)=\sum_{i\leq n}\left[\left(W_{Q}x_{n}\right)^{\intercal}\left(W_{K}x_{i}\right)+B_{i,n}\right]W_{V}x_{i},$ (2)
where $W_{Q},W_{K},W_{V}$ are the query, key, value matrices, respectively, and $B\in{\mathbb{R}}^{N\times N}$ is the position bias. To focus on the impact of the position embeddings, we fix $W_{V}=I$ and consider position-only attention score. Then we have
$f_{\text{PLA}}\left(x;B\right)=\sum_{i\leq n}B_{i,n}x_{i}=\sum_{i\leq n}e_{i}^{\intercal}Be_{n}x_{i}.$
Rewriting $B$ as $A$ and restricting $A$ to an upper triangular matrix, we have
$f_{\text{PLA}}\left(x;A\right)=\sum_{i=1}^{n}e_{i}^{\intercal}Ae_{n}x_{i}=x^{\intercal}Ae_{n},\quad\text{ where }A\in{\mathcal{U}}_{N}.$
In the computation of attention, the query position is asymmetric to other positions and can reflect the scale of the instance. To formalize the intuition, we introduce the notation $n(x):=\operatorname*{arg\,max}_{i\in[N]}\{x_{i}=-1\}$, which represents the last dimension where $x_{i}=-1$. In the LDHD generalization framework, $n(x)$ can be interpreted as the lowest dimension of the subspaces containing $x$, reflecting the dimension of $x$. We take $n(x)$ as the query position of $x$. Specifically, when $x$ is all-ones, it is treated as an ”empty” sequence. In this case, we fix its output to align with the definition of PLAA. By the above derivation, we obtain Definition 5.
B.2 Construction of PLAA with APE
The expression of the linear attention with APE is slightly different from (2). In practice, APE is commonly added to the token embeddings. Therefore, we omit the position bias and add APE to each $x_{i}$ when computing the attention score (we slightly abuse the notation of $x_{i}$ to denote the token embedding), i.e.,
$f_{\text{LA}}^{\text{APE}}\left(x;W_{Q},W_{K},W_{V},\left\{p_{i}\right\}_{i\in[N]}\right)=\sum_{i\leq n}\left[W_{Q}\left(x_{n}+p_{n}\right)\right]^{\intercal}\left[W_{K}\left(x_{i}+p_{i}\right)\right]W_{V}x_{i},$
where $p_{i}$ is the position embedding for the position $i$.
Similar to the derivation in Section B.1, we have
$f_{\text{PLAA}}^{\text{APE}}\left(x;W_{Q},W_{K}\right)=\sum_{i\leq n(x)}p_{n(x)}^{\intercal}W_{Q}^{\intercal}W_{K}p_{i}x_{i}=\sum_{i\leq n(x)}e_{n(x)}^{\intercal}P^{\intercal}W_{Q}^{\intercal}W_{K}Pe_{i}x_{i},$
where $P$ is the learnable position embedding matrix such that $p_{i}=Pe_{i}$ for all $i\in[N]$. Without loss of generality, we absorb the learnable parameters $W_{Q},W_{K}$ into $P$, obtaining
$f_{\text{PLAA}}^{\text{APE}}\left(x;P\right)=\sum_{i\leq n}x_{i}^{\intercal}e_{i}^{\intercal}P^{\intercal}Pe_{n(x)}=x^{\intercal}\left(M_{N}^{u}\circ P^{\intercal}P\right)e_{n(x)}.$
B.3 Construction of PLAA with GRPE
RPE can be seen as a reparameterization of the position bias matrix such that the entries on the same diagonals share learnable parameters, i.e.,
$B=\sum_{i=1}^{N}D_{i}p_{i},$
where $D_{k}$ is the $k$-th upper diagonal matrix and $p_{i}$ is the learnable parameter for all $1\leq k\leq N$.
We generalize RPE so that our model can cover more general position biases. A natural extension is to choose general upper triangular matrices, denoted by ${\mathcal{U}}_{\text{GRPE}}=\{U_{1},\dots,U_{r}\}$, instead of ${\mathcal{U}}_{\text{RPE}}=\{D_{1},\dots,D_{N}\}$. To normalize, we suppose $\langle U_{i},U_{i}\rangle=1$ for all $i=1,\dots,r$. (This condition does not hold for ${\mathcal{U}}_{\text{RPE}}$. We can slightly modify the matrices by replacing $D_{k}$ with $\frac{1}{N+1-k}D_{k}$ to satisfy the condition.) We further require $(U_{i})_{kl}(U_{j})_{kl}=0$ for all $i\neq j$ and $1\leq k,l\leq N$, i.e., each $B_{i,j}$ is characterized by at most one learnable parameter. Using the reparameterization
$B=\sum_{i=1}^{r}U_{i}p_{i}$
in the derivation of PLAA, we obtain PLAA with GRPE in Definition 7.
Appendix C Proofs
C.1 Proof for Theorem 1
@@
$V_{1}(c)=\sum_{f\in{\mathcal{F}}/\left(c|_{{\mathcal{X}}_{N_{0}}}\right)}\mathbb{E}_{x\sim{\mathcal{D}}_{1}}\left[\ell\left(c(x),f(x)\right)\right],$
$V_{2}(c)=\sum_{f\in{\mathcal{F}}/\left(c|_{{\mathcal{X}}_{N_{0}}}\right)}\mathbb{E}_{x\sim{\mathcal{D}}_{2}}\left[\ell\left(c(x),f(x)\right)\right].$
@@
$\overset{(a)}{=}\sum_{f\in{\mathcal{F}}/\left(c_{2}|_{{\mathcal{X}}_{N_{0}}}\right)}\mathbb{E}_{x\sim{\mathcal{D}}_{1}}\left[\ell\left(c_{2}(x),c_{2}(x)\right)\right]$
@@
By the no-free-lunch theorem (Wolpert & Macready, 1997; Wolpert, 2002) for ${\mathcal{X}}_{N}\setminus{\mathcal{X}}_{N_{0}}$, $\mathcal{Y}$}, and ${\mathcal{F}}_{{\mathcal{X}}_{N}\setminus{\mathcal{X}}_{N_{0}},{\mathcal{Y}}}$, we have
@@ C.2 Proof for Theorem 2
According to the proof of Theorem 3.8 in Abbe et al. (2023), the RFM model $f_{\text{RFM}}(z;a)$ converges to the min-degree interpolator (w.r.t. the variable $z$) in ${\mathcal{G}}^{\prime}_{N_{0},c}$, i.e.,
@@
C.3 Proof for Corollary 1
Lemma 1.
For any $f\in{\mathcal{F}}$, the min-degree interpolator $f^{*}$ on ${\mathcal{X}}_{N_{0}}$ satisfies $I(f^{*})\subset[N_{0}]$.
Proof for Lemma 1.
Assume that there exist some $f\in{\mathcal{F}}$ such that the min-degreee interpolator $\hat{f}$ of $f$ on ${\mathcal{X}}_{N_{0}}$ does not satisfy $I(\hat{f})\subset[N_{0}]$.
Let $\tilde{f}(x)$ be the function constructed by fixing all $x_{i},i\not\in[N_{0}]$ to 1 in $\hat{f}(x)$. By the construction, we have $\tilde{f}(x)=\hat{f}(x)$ for all $x\in{\mathcal{X}}_{N_{0}}$ and thus $\tilde{f}(x)$ is also an interpolator of $f$ on ${\mathcal{X}}_{N_{0}}$.
[…]
By Lemma 1, for any $f\in{\mathcal{F}}$ such that $I(f)\not\subset[N_{0}]$, the min-degreee interpolator $\hat{f}$ is not identical to $f$ on ${\mathcal{X}}_{N}$ and thus does not achieves LDHD generalization from ${\mathcal{X}}_{N_{0}}$ to ${\mathcal{X}}_{N}$.
C.4 Proof for Theorem 3
Let $U({\mathcal{X}}_{N_{0}})$ be the uniform distribution over ${\mathcal{X}}_{N_{0}}$. The loss function is
$L(P)$ $=\mathbb{E}_{x\sim U\left({\mathcal{X}}_{N_{0}}\right)}\left[\frac{1}{2}\left(f_{\text{PLAA}}^{\text{APE}}(x;P)-f_{\text{PLAA}}(x;A^{*})\right)^{2}\right]$
$=\mathbb{E}_{x\sim U\left({\mathcal{X}}_{N_{0}}\right)}\left[\frac{1}{2}\left(x^{\intercal}\left(M_{N}^{u}\circ P^{\intercal}P\right)e_{n(x)}-x^{\intercal}A^{*}e_{n(x)}\right)^{2}\right]$
$=\mathbb{E}_{x\sim U\left({\mathcal{X}}_{N_{0}}\right)}\left[\frac{1}{2}\left(\left\langle xe_{n(x)}^{\intercal},M_{N}^{u}\circ P^{\intercal}P\right\rangle-\left\langle xe_{n(x)}^{\intercal},A^{*}\right\rangle\right)^{2}\right]$
$\overset{(a)}{=}\mathbb{E}_{x\sim U\left({\mathcal{X}}_{N_{0}}\right)}\left[\frac{1}{2}\left(\left\langle xe_{n(x)}^{\intercal},M_{N}^{u}\circ P^{\intercal}P\right\rangle-\left\langle xe_{n(x)}^{\intercal},M_{N}^{u}A^{*}\right\rangle\right)^{2}\right]$
$=\mathbb{E}_{x\sim U\left({\mathcal{X}}_{N_{0}}\right)}\left[\frac{1}{2}\left(\left\langle M_{N}^{u}\circ xe_{n(x)}^{\intercal},P^{\intercal}P-A^{*}\right\rangle\right)^{2}\right],$
where (a) follows from the fact that $A^{*}$ is upper triangular.
Lemma 2.
For any matrix $Z\in{\mathbb{R}}^{N\times N}$, it holds that
$\mathbb{E}_{x\sim U\left({\mathcal{X}}_{N_{0}}\right)}\left[\frac{1}{2}\left(\left\langle M_{N}^{u}\circ xe_{n(x)}^{\intercal},Z\right\rangle\right)^{2}\right]=\frac{1}{2}\left\|Q_{N_{0}}\circ Z\right\|_{F}^{2},$
where
$Q_{N_{0}}=\begin{bmatrix}2^{-\frac{N_{0}}{2}}&\cdots&\cdots&2^{-\frac{1}{2}}&0&\cdots&0\\ 0&2^{-\frac{N_{0}-1}{2}}&\cdots&2^{-\frac{1}{2}}&0&\cdots&0\\ \vdots&\ddots&\ddots&2^{-\frac{1}{2}}&0&\cdots&0\\ 0&\cdots&0&2^{-\frac{1}{2}}&0&\cdots&0\\ 0&\cdots&\cdots&0&0&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&0&\cdots&0\\ 0&\cdots&0&0&0&\cdots&0\\ \end{bmatrix}\in{\mathbb{R}}^{N\times N},$
that is,
$\left(Q_{N_{0}}\right)_{ij}=\begin{cases}2^{\frac{N_{0}-j+1}{2}},&1\leq i\leq j\leq N_{0},\\ 0,&\text{otherwise}.\end{cases}$
Proof for Lemma 2.
Note that
$\mathbb{E}_{x\sim U\left({\mathcal{X}}_{N_{0}}\right)}\left[\frac{1}{2}\left(\left\langle M_{N}^{u}\circ xe_{n(x)}^{\intercal},Z\right\rangle\right)^{2}\right]=$ $\frac{1}{2^{N_{0}}}\sum_{k=1}^{N_{0}}\sum_{x:n(x)=k}\frac{1}{2}\left(\left\langle M_{N}^{u}\circ xe_{k}^{\intercal},Z\right\rangle\right)^{2}$
$=$ $\frac{1}{2^{N_{0}}}\sum_{k=1}^{N_{0}}\sum_{x:n(x)=k}\frac{1}{2}\left(\sum_{i=1}^{k-1}x_{i}Z_{ik}-Z_{kk}\right)^{2}$
$=$ $\frac{1}{2^{N_{0}}}\sum_{k=1}^{N_{0}}\sum_{x:n(x)=k}\frac{1}{2}\sum_{i=1}^{k}Z_{ik}^{2}$
$\overset{(a)}{=}$ $\frac{1}{2^{N_{0}}}\sum_{k=1}^{N_{0}}\frac{2^{k-1}}{2}\sum_{i=1}^{k}Z_{ik}^{2}$
$=$ $\frac{1}{2}\sum_{k=1}^{N_{0}}\sum_{i=1}^{k}\left(2^{-\frac{N_{0}-k+1}{2}}Z_{ik}\right)^{2}=\frac{1}{2}\left\|Q_{N_{0}}\circ Z\right\|_{F}^{2},$
where (a) follows from the fact that $\left|\{x\mid n(x)=k\}\right|=2^{k-1}$. ∎
According to Lemma 2, we have
$L(P)=\frac{1}{2}\left\|Q_{N_{0}}\circ\left(P^{\intercal}P-A^{*}\right)\right\|_{F}^{2}.$ (3)
For any $0<\alpha\leq\alpha_{0}$, we have $L\left(P_{\infty}(\alpha)\right)=0$ and thus
$\left(M_{N}^{u}\circ P_{\infty}(\alpha)^{\intercal}P_{\infty}(\alpha)\right)_{[N_{0}],[N_{0}]}=A^{*}_{[N_{0}],[N_{0}]}.$ (4)
Let $p_{i}\in{\mathbb{R}}^{d_{P}}$ denote the $i$-th column of $P$. According to (3), we have
$\frac{\partial L}{\partial p_{i}}=0,$
for all $i=N_{0}+1,\dots,N$. Therefore, $\left(p_{i}\right)_{t}=\left(p_{i}\right)_{0}$ for all $t$ and $i=N_{0}+1,\dots,N$.
Define
$\tilde{A}^{*}:=\begin{cases}\left(A^{*}\right)_{ij},&1\leq i\leq j\leq N_{0},\\ 0,&\text{otherwise}.\end{cases}$
Let $p^{\alpha}_{i}$ be the $i$-th column of $P_{\infty}(\alpha)$. We have
$\left\|M_{N}^{u}\circ P_{\infty}^{\intercal}(\alpha)P_{\infty}(\alpha)-\tilde{A}^{*}\right\|_{F}^{2}=$ $\sum_{1\leq i\leq j\leq N}\left(\left(p_{i}^{\alpha}\right)^{\intercal}p_{j}^{\alpha}-\tilde{A}^{*}_{ij}\right)^{2}$
$\overset{(a)}{=}$ $\sum_{j=N_{0}+1}^{N}\sum_{i=1}^{j}\left(\left(p_{i}^{\alpha}\right)^{\intercal}p_{j}^{\alpha}\right)^{2}$
$\leq$ $\sum_{j=N_{0}+1}^{N}\sum_{i=1}^{j}\left\|p_{i}^{\alpha}\right\|_{2}^{2}\left\|p_{j}^{\alpha}\right\|_{2}^{2}$
$=$ $\sum_{j=N_{0}+1}^{N}\sum_{i=1}^{N_{0}}\left\|p_{i}^{\alpha}\right\|_{2}^{2}\left\|p_{j}^{\alpha}\right\|_{2}^{2}+\sum_{j=N_{0}+1}^{N}\sum_{i=N_{0}+1}^{j}\left\|p_{i}^{\alpha}\right\|_{2}^{2}\left\|p_{j}^{\alpha}\right\|_{2}^{2}.$
where (a) follows from (4) and the definition of $\tilde{A}^{*}$.
Note that $\left\|p_{i}^{\alpha}\right\|_{2}^{2}=A_{ii}^{*}\leq\left\|A^{*}\right\|_{\infty}$ for all $1\leq i\leq N_{0}$ and $\left\|p_{i}^{\alpha}\right\|_{2}^{2}=\alpha$ for all $N_{0}+1\leq i\leq N$. Then we have
$0\leq\left\|M_{N}^{u}\circ P_{\infty}^{\intercal}(\alpha)P_{\infty}(\alpha)-\tilde{A}^{*}\right\|_{F}^{2}\leq N(N-N_{0})\alpha\left\|A^{*}\right\|_{\infty}+\frac{(N-N_{0}+1)(N-N_{0})}{2}\alpha^{2}.$
As $\alpha\to 0$, we have
$0\leq\left\|M_{N}^{u}\circ\hat{P}^{\intercal}\hat{P}-\tilde{A}^{*}\right\|_{F}^{2}\leq 0,$
which means $M_{N}^{u}\circ\hat{P}^{\intercal}\hat{P}=\tilde{A}^{*}$.
[…]
$b_{ij}^{\text{PLAA}}(x)=\left\langle xe_{n(x)}^{\intercal},E_{ij}\right\rangle.$
Then we have
$f_{\text{PLAA}}^{\text{APE}}(x;\hat{P})=\sum_{1\leq i\leq j\leq N}\tilde{A}^{*}_{ij}b_{ij}^{\text{PLAA}}(x)=\sum_{1\leq i\leq j\leq N_{0}}A^{*}_{ij}b_{ij}^{\text{PLAA}}(x).$
For any $g=\sum_{1\leq i\leq j\leq N}A_{ij}b_{ij}^{\text{PLAA}}(x)\in{\mathcal{G}}_{N_{0},A^{*}}^{\text{PLAA}}$, we have $A_{ij}=A^{*}_{ij}$ for all $1\leq i\leq j\leq N_{0}$. Since $A_{ij}^{2}\geq\tilde{A}_{ij}^{2}$ for all $1\leq i\leq j\leq N$, we have
$\operatorname{DegP}_{{\mathcal{B}}_{N}^{\text{PLAA}}}\left(f_{\text{PLAA}}^{\text{APE}}(x;\hat{P})\right)\leq\operatorname{DegP}_{{\mathcal{B}}_{N}^{\text{PLAA}}}\left(g\right).$
[…]
C.5 Proof for Theorem 4